# Quench dynamics via recursion method and dynamical quantum phase transitions

### v1

@@ Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis, which is recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. In this work, we extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension we find a clear evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
@@
Introduction.  Addressing quantum dynamics by means of the recursion method has a long history Mori (1965); Dupuis (1967); Haydock (1980); Mattis (1981); Joslin and Gray (1986); Roldan et al. (1986); Viswanath and Müller (2008); Lindner and Auerbach (2010); Nandy et al. (2024). In simple words, the (Heisenberg version of the) method amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about. A new impact has been given to the method by the universal operator growth hypothesis Parker et al. (2019), which has informed about the generic pattern of the evolution operator in the Lanczos basis.
@@
Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively by computing infinite-temperature autocorrelation functions and transport coefficients Parker et al. (2019); Khait et al. (2016); de Souza et al. (2020); Yates et al. (2020a, b); Yuan et al. (2021); Wang et al. (2024); Bartsch et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024). There the method has demonstrated high efficiency. This efficiency is particularly notable in application to two-dimensional systems Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024), where traditional numerical methods often struggle. Another notable advantage of the recursion method is that it allows a straightforward symbolic implementation, when a single computation covers all parameters of the Hamiltonian Uskov and Lychkovskiy (2024); Teretenkov et al. (2024).
@@
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2024); Loizeau et al. (2025a). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to try to apply it to quench dynamics.
@@
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition. Then some results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions posed by them.
Recursion method for quench dynamics.  Consider an observable given by a self-adjoint Shrödinger operator $A$. The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system. The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian. $A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$. Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$. Here a scalar product in the space of operators is introduced according to $\big{(}A|B\big{)}\equiv\operatorname{\mbox{tr}}\big{(}A^{\dagger}B\big{)}/d$, where $d$ is the Hilbert space dimension (which is assumed to be finite). The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$. The coefficients $b_{n}$ are referred to as Lanczos coefficients. In this basis the Liouvillian has the tridiagonal form,
@@
$A_{t}=\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
where $\phi^{n}_{t}\equiv(O^{n}|A_{t})$ are the solution of the system of coupled equations, which can be written in the matrix form as
@@
In practice, one can explicitly compute only a finite number $n_{\rm max}$ of Lanczos coefficients. The rest of the coefficients should be extrapolated. The universal operator growth hypothesis Parker et al. (2019) (see also precursor work Liu and Müller (1990); Florencio et al. (1992); Zobov and Lundin (2006); Elsayed et al. (2014); Bouch (2015)) states that in a generic quantum many-body system Lanczos coefficients $b_{n}$ scale asymptotically linearly with $n$ at large $n$ (with a logarithmic correction for one-dimensional systems). This asymptotic behavior has been confirmed in various many-body models Cao (2021); Noh (2021); Heveling et al. (2022); Uskov and Lychkovskiy (2024); De et al. (2024). As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients. However, extrapolating Lanczos coefficients by a linear function is a too rough approximation: the qualitative behavior of the solutions of the Heisenberg equation strongly depends on the subleading asymptotic terms Viswanath et al. (1994); Parker et al. (2019); Yates et al. (2020a, b); Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Uskov and Lychkovskiy (2024); Dodelson (2025). Guided by the results of ref. Gam , we employ the following extrapolation formula:
@@
The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gam , which is generally the case. Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
@@
Note that $\|A\|\,\phi^{0}_{t}=d^{-1}\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$. Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al. (2019); Wang et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024). The out-of-equilibrium post-quench expectation value $\langle A_{t}\rangle$ is, however, a more complex object. Below we explore it in specific models.
@@
Figure 2: Scaling of the estimated breakdown time $t_{*}$ with respect to the number $n_{\rm max}$ of explicitly know Lanczos coefficients. Circles, squares and diamonds correspond to pure initial states polarized in $x$, $y$ and $z$ direction, respectively. The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition. Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
@@
We are able to compute $n_{\rm max}=45$ nested commutators and Lanczos coefficients. The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method is presented in Fig. 1. It is benchmarked by the results of exact diagonalization (performed with QuSpin Weinberg and Bukov (2017)) and time-dependent variational principle empowered by matrix product states (performed with ITensor Fishman et al. (2022a, b)). In one dimension the latter methods are essentially exact (and, of course, produce the same result) up to large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ at initial time, but then abruptly breaks down. We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit. Namely, the breakdown time $t_{*}$ is very close tho the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig. 1.
The breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient. However, this growth is saturated at $t_{c}$, as shown in Fig. 2. This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig. 3 (a). Note that the one-dimensional Ising model is known for its atypical thermalization properties Bañuls et al. (2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024), including long-lived Lin and Motrunich (2017) or even eternal Bañuls et al. (2011) oscillations of polarization around its time-averaged value seen in Fig. 3 (a).
@@
Figure 3: Quench dynamics on (a) one-dimensional nearest-neighbor, (b) square, and (c) cubic lattices. Shown is the expectation value of the total magnetization in the $z$-direction. The initial state is the pure product state (9) polarized along the same direction. Blue solid line is the result of the recursion method. The chosen time span is sufficient for the latter to equilibrate. (a) One-dimensional Ising model (7) with $h_{x}=h_{z}=1$ (the same as in Fig. 1 but for a larger time span). Orange is the exact numerical result. The horizontal dashdotted line is the numerically time-averaged post-quench value of the observable. (b) and (c): The transverse-filed Ising model (10) with $h_{z}=1$ on the square and cubic lattices, respectively. Points are the results of the sparse Pauli dynamics method Begušić and Chan (2024).
@@
where ${\mathbf{i}}$ and ${\mathbf{j}}$ enumerate lattice sites and the first sum runs over pairs of neighbouring sites.
We compute $22$ and $12$ nested commutators, respectively. Some results are presented in Fig. 3 (b),(c). They are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2024); Begušić .
@@
Conclusions.  In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states. We have applied the method to quantum spin systems in one, two, and three dimensions. In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down. Notably, we found that this breakdown time closely coincides with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question. The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce. We have compared our results with those obtained from the sparse Pauli dynamics method Begušić and Chan (2024); Begušić and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension. However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs. Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an open question.
@@
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025b). The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method. It does not report neither the finite convergence domain of the method nor an association with the DQPT.

### v2

@@ Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. We extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension we find an evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
@@
Introduction.  Addressing quantum dynamics by means of the recursion method has a long history Mori (1965); Dupuis (1967); Haydock (1980); Mattis (1981); Joslin and Gray (1986); Roldan et al. (1986); Viswanath and Müller (2008); Lindner and Auerbach (2010); Nandy et al. (2024). The (Heisenberg version of the) method essentially amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about. A new impact has been given to the method by the universal operator growth hypothesis Parker et al. (2019) – a conjecture about a generic form of the evolution generator in the Lanczos basis.
@@
Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively to computing infinite-temperature autocorrelation functions and transport coefficients Parker et al. (2019); Khait et al. (2016); de Souza et al. (2020); Yates et al. (2020a, b); Yuan et al. (2021); Wang et al. (2024); Bartsch et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024); Füllgraf et al. (2025). There the method has demonstrated high efficiency. This efficiency is particularly notable in application to two-dimensional systems Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024), where traditional numerical methods often struggle. Another notable advantage of the recursion method is that it allows a straightforward symbolic implementation, when a single computation covers all parameters of the Hamiltonian Uskov and Lychkovskiy (2024); Teretenkov et al. (2024).
@@
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al. (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2025); Loizeau et al. (2025a, b). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
@@
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition. Then results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions raised.
[…]
@@
$A_{t}=\|A\|\,\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
where $\phi^{n}_{t}\equiv(O^{n}|A_{t})/\|A\|$ are the solution of the system of coupled equations, which can be written in the matrix form as
@@
In practice, one can explicitly compute only a finite number $n_{\rm max}$ of Lanczos coefficients. The rest of the coefficients should be extrapolated. The universal operator growth hypothesis Parker et al. (2019) (see also precursor work Liu and Müller (1990); Florencio et al. (1992); Zobov and Lundin (2006); Elsayed et al. (2014); Bouch (2015)) states that in a generic quantum many-body system Lanczos coefficients $b_{n}$ scale asymptotically linearly with $n$ at large $n$ (with a logarithmic correction for one-dimensional systems). This asymptotic behavior has been confirmed in various many-body models Cao (2021); Noh (2021); Heveling et al. (2022); Uskov and Lychkovskiy (2024); De et al. (2024); Ermakov (2025). As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients. However, extrapolating Lanczos coefficients by a linear function is a too rough approximation: the qualitative behavior of the solutions of the Heisenberg equation strongly depends on the subleading asymptotic terms Viswanath et al. (1994); Parker et al. (2019); Yates et al. (2020a, b); Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Uskov and Lychkovskiy (2024); Dodelson (2025); Gamayun et al. (2025). Guided by the results of ref. Gamayun et al. (2025), we employ the following extrapolation formula:
@@
The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gamayun et al. (2025), which is generally the case. Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
@@
It should be emphasized that the sequence $\langle O^{n}\rangle$ generally does not have any apparent structure and is not expected to have any particular asymptotic behavior, in contrast to the sequence of Lanczos coefficients $b_{n}$. We further discuss this fact and illustrate it by examples in the Supplement sup .
Note that $\phi^{0}_{t}=({\cal N}/\|A\|)\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$, up to a constant factor. Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al. (2019); Wang et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024). The out-of-equilibrium post-quench expectation value $\langle A_{t}\rangle$ is, however, a more complex object. Below we explore it in specific models.
@@
Figure 2: Scaling of the estimated breakdown time $t_{*}$ with the number $n_{\rm max}$ of explicitly know Lanczos coefficients for various initial polarizations. $t_{*}$ is defined as the time when the recursion method result departs from the essentially exact numerical result by 0.005. The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition, see Fig. 1 (a). Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
@@
We are able to compute $n_{\rm max}=48$ nested commutators and Lanczos coefficients for the Hamiltonian (7) and the observable (8). The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method, i.e. from eq. (6), is presented in Fig. 1. Values of $b_{n}$, as well as individual ingredients entering eq. (6), i.e. $\phi^{n}_{t}$ and $\langle O^{n}\rangle$, are presented in the Supplement sup .
@@
Figure 3: Quench dynamics of nonintegrable Ising models on (a) one-dimensional, (b) square, and (c) cubic lattices. Shown is the expectation value of the total magnetization in the $z$-direction. The initial state is the pure product state (9) polarized along the same $z$ direction. Blue solid line is the result of the recursion method. The chosen time span is sufficient for the latter to equilibrate. (a) One-dimensional Ising model (7) with $h_{x}=h_{z}=1$ (the same as in Fig. 1 but for a larger time span). Orange is the exact numerical result. The horizontal dashdotted line is the numerically time-averaged post-quench value of the observable. (b) and (c): The transverse-filed Ising model (10) with $h_{z}=1$ on the square and cubic lattices, respectively. Points are the results of the sparse Pauli dynamics method Begušić and Chan (2025). Insets show explicitly computed Lanczos coefficients (points) and their extrapolation according to eqs. (4), (5) (black lines).
The results of the recursion method are benchmarked by the results of exact diagonalization performed with QuSpin Weinberg and Bukov (2017) and time-dependent variational principle empowered by matrix product states performed with ITensor Fishman et al. (2022a, b) (see the Supplement sup for details). In one dimension these numerical methods are essentially exact up to sufficiently large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down. We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit.111 There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al. (2006); Eckstein et al. (2009); Uhlmann et al. (2010); Garrahan and Lesanovsky (2010); Diehl et al. (2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions. In contrast, the definition we adopt was introduced later Heyl et al. (2013) and is based on a formal analogy between the Loschmidt rate and the free energy. The breakdown time $t_{*}$ is very close to the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig. 1.
More precisely, the breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient, but this growth saturates at $t_{c}$, as shown in Fig. 2. This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig. 3 (a). Note that the one-dimensional transverse-field Ising model is known for its unusually slow thermalization Bañuls et al. (2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024) which is seen in Fig. 3 (a).
@@
where ${\mathbf{i}}$ and ${\mathbf{j}}$ enumerate lattice sites and the first sum runs over pairs of neighbouring sites. The observable is the same as in the previous case – the total magnetization in the $z$-direction, $A=\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}}$.
We compute $23$ and $12$ nested commutators, respectively. Time evolution of the magnetization is presented in Fig. 3 (b),(c). The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2025); Beg .
@@
Conclusions and discussion.  In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states. We have applied the method to quantum spin systems in one, two, and three dimensions.
In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down.
The breakdown indicates that the expansion (6) of a Heisenberg observable in the Lanczos basis has a finite convergence domain, at least for one-dimensional systems. This fact may limit the scope of concepts and techniques based on this expansion Parker et al. (2019); Chu et al. (2024); Loizeau et al. (2025c).
Notably, the breakdown time turns out to closely coincide with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension. The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question. The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce. We have compared our results to those obtained from the sparse Pauli dynamics method Begušić and Chan (2025); Beg and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension. However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs. Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an important open question.
@@
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025c). The latter paper also addresses quench dynamics in one-dimensional spin systems by means of the recursion method. It does not report the finite convergence domain of the method or any association with the DQPT.